OpenStudy (anonymous):

solve: e^integral(-2/t)dt Any help is appreciated. Kind regards

OpenStudy (lgbasallote):

\[\huge e^{\int -\frac 2t dt}\] do you know the value of \[\huge \int -\frac 2t dt\]

OpenStudy (anonymous):

I think it is -2ln(t)

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

now apply the log rule.. \[\huge a \ln b \implies \ln b^a\]

OpenStudy (lgbasallote):

so what will be the value of -2 ln t

OpenStudy (anonymous):

-1 minute plse ...screen has gone blank

OpenStudy (lgbasallote):

sure sure

OpenStudy (anonymous):

busy powering up imac ...1 more minute plse

OpenStudy (lgbasallote):

no rush :D

OpenStudy (anonymous):

I'm assuming you want the integral of e^(2t) dt The integral is simply (1/2)*e^(2t) + c (c is a constant) This is because the integral of an exponential, e^(at) dt is: (1/a)e^(at) + c

OpenStudy (anonymous):

for some reason openstudy wont open on imac...this could be a while ...sorry....let me log off on the windos pc...mebbe that the problem ;)

OpenStudy (lgbasallote):

haha take your time. just tag me when you're ready

OpenStudy (anonymous):

the answer is lnt^-2

OpenStudy (lgbasallote):

right.. so.. \[\LARGE e^{\int -\frac 2t dt} \implies e^{-2 \ln t} \implies e^{\ln (t^{-2})}\]

OpenStudy (anonymous):

okay

OpenStudy (lgbasallote):

do you know how to simplify that?

OpenStudy (anonymous):

yes, t^-2 = 1/t^2

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

does the ln cancel the e out

OpenStudy (lgbasallote):

yup \[\huge e^{\ln a} = a\]

OpenStudy (anonymous):

excellent - where do i find all these rules..do you know of a good resource

OpenStudy (lgbasallote):

i have written a list of log rules before in this site...it also comes with a brief description and examples... http://openstudy.com/updates/4fa45f8fe4b029e9dc34e0b5

OpenStudy (anonymous):

Wow...That's amazing....I didn't even know we had such excellent resources here on openstudy....Thank you very much ....Much appreciated

OpenStudy (lgbasallote):

welcome ^_^ btw...you wouldnt be happening to be studying integrating factors in differential equations, would you?

OpenStudy (anonymous):

yep...thats exactly why i posted this question..it is part of an Integrating Factor problem...How did you guess.?

OpenStudy (lgbasallote):

because it didnt look like a normal integration problem =)) the only time i saw that kind of integration is from integrating factor

OpenStudy (anonymous):

You are spot on....nice! Its not too hard now...we move on to 2nd order next semester. Is that difficult?

OpenStudy (lgbasallote):

well...hmm...not really...you just have to remember solutions

OpenStudy (anonymous):

okay...If you have any more tips on how to solve these IF problems that would be much appreciated. I struggle sometimes with the exp function...but with the help of the great people (like you) at openstudy my task has been made more easier... :)

OpenStudy (lgbasallote):

hmm i dont know shortcuts to IF haha

OpenStudy (anonymous):

okay...;) anyway ..thanks for your help .....i really aprreciate it...take care . :)

OpenStudy (lgbasallote):

welcome ^_^